Shen is putting money into a savings account. He starts with 650 in the savings account, and each week he adds $30.
Let S represent the total amount of money in the savings account (in dollars), and let W represent the number of weeks Shen has been adding money. Write an equation relating S to W. Then use this equation to find the total amount of money in the savings account after 11 weeks.

Answer by math_tutor2020(3817)


W = number of weeks
S = total amount of money in the savings account (in dollars)

Shen starts off with $650.
After 1 week, he adds on $30
The balance would be 650+30 which I'll leave unsimplified.

After 2 weeks, he adds on another 30, so,
balance = (650+30)+30 = 650+(30+30) = 650+2*30

Then after 3 weeks, his balance is...
balance = (650+2*30)+30 = 650+(2*30+30) = 650+3*30

At this point you can probably spot the pattern
balance = 650+(number of weeks)*30
which simplifies to
S = 650+W*30
S = 30W+650

Now plug in W = 11 to find how much is in the account after 11 weeks.
S = 30W+650
S = 30*11+650
S = 330+650
S = 980

----------------------------------------------------

Answers:

The equation is S = 30W+650
After 11 weeks, Shen has $980 in the account.